THE WIDTH OF A RECTANGULAR PLAYGROUND IS 2x-5 FEET, AND THE LENGTH IS 3x+9 FEET. WRITE A POLYNOMINAL P(x) THAT REPRESENTS THE PERIMETER THEN EVALUATE THIS OERAMETER IF x IS 4 FEET.
Answer by stanbon(75887)

THE WIDTH OF A RECTANGULAR PLAYGROUND IS 2x-5 FEET, AND THE LENGTH IS 3x+9 FEET. WRITE A POLYNOMINAL P(x) THAT REPRESENTS THE PERIMETER THEN EVALUATE THIS OERAMETER IF x IS 4 FEET.
------------------------------------
Perimeter = 2(width + length)
P = 2(2x-5+3x+9)
P = 2(5x+4)
----------------
P(4) = 2(24) = 48 feet
========================
Cheers,
Stan H.